OpenStudy (anonymous):

Find an exact solution for r5-1/x-r5/2 . Then, find the approximate solution. A. 10-2r5/5;1.1 B. 2-2rs;-2.5 C. -2 ; 2 D. 2r5-2/r5;1.1

OpenStudy (anonymous):

r is for sqaure root

OpenStudy (anonymous):

someone please help

OpenStudy (unklerhaukus):

\[Q:\qquad\sqrt5-\frac1x-\frac{\sqrt5}2\] \[A.\qquad \frac{10-2\sqrt5}5;\qquad1.1\]\[ B.\qquad 2-2\sqrt 5;\qquad-2.5\]\[C.\qquad -2 ;\qquad\qquad \qquad 2 \]\[D. \qquad 2\sqrt5-\frac{2}{\sqrt 5};\qquad 1.1\]

OpenStudy (unklerhaukus):

what is x

OpenStudy (anonymous):

it is what you have to find

OpenStudy (unklerhaukus):

i think something is missing, maybe =0

OpenStudy (anonymous):

maybe

OpenStudy (unklerhaukus):

is this you question \[Q:\qquad\sqrt5-\frac1x-\frac{\sqrt5}2=0\]

OpenStudy (anonymous):

i need to find the exact solution

OpenStudy (unklerhaukus):

is the equal zero part in the question, because im just guessing

OpenStudy (anonymous):

I guess not sure we can use 0

OpenStudy (turingtest):

@chadandtricia_2009 please be sure to type out your question in full use parentheses when needed to avoid confusion

OpenStudy (unklerhaukus):

\[\sqrt5-\frac1x-\frac{\sqrt5}2=\frac{\sqrt5}2-\frac1x\]

OpenStudy (turingtest):

I was thinking that\[{\sqrt5-1\over x}-\frac{\sqrt5}2=0\]was the problem most likely @chadandtricia_2009 please clarify

OpenStudy (anonymous):

all I have is the equation no equals

OpenStudy (turingtest):

then it's not an equation by definition

OpenStudy (turingtest):

equations have to have = signs without an = sign it is not an equation, and cannot be solved for x

OpenStudy (anonymous):

just need to find the exact solution for the problem

OpenStudy (turingtest):

that is impossible without an = sign you could simplify the expression a little perhaps, but you cannot solve for x

OpenStudy (unklerhaukus):

C. is definitely not right \[-2\cancel \approx 2\]

OpenStudy (turingtest):

so look for a typo, there is one somewhere

OpenStudy (anonymous):

I'm just going to guess at it